# Entanglement Rényi Negativity of Interacting Fermions from Quantum Monte Carlo Simulations

### v2

@@ Abstract
Many-body entanglement unveils additional aspects of quantum matter and offers insights into strongly correlated physics. While ground-state entanglement has received much attention in the past decade, the study of mixed-state quantum entanglement using negativity in interacting fermionic systems remains largely unexplored. We demonstrate that the partially transposed density matrix of interacting fermions, similar to their reduced density matrix, can be expressed as a weighted sum of Gaussian states describing free fermions, enabling the calculation of rank-$n$ Rényi negativity within the determinant quantum Monte Carlo framework. We conduct the first calculation of rank-two Rényi negativity for the half-filled Hubbard model and the spinless $t$-$V$ model and find that the area law coefficient of the Rényi negativity has a logarithmic finite-size scaling at the finite-temperature transition point. Our work contributes to the calculation of entanglement and sets the stage for future studies on quantum entanglement in various fermionic many-body mixed states.
@@
Introduction.—The characterization of emerging quantum many-body phenomena is multifaceted. Traditionally, physicists have relied on local measurements based on linear response to investigate matter. In recent decades, the utilization of quantum entanglement, a fundamental concept in quantum physics and a powerful tool in quantum information, has become pivotal in unveiling the additional aspects of quantum matter, including the identification of exotic phases and quantum criticality [1, 2, 3]. A prominent example is the entanglement entropy (EE) used in bipartite ground-state entanglement studies [4], where various corrections to the area law [5, 6, 7] have been employed to classify quantum phases. These include logarithmic corrections in the leading area-law term for 1D critical systems [8, 9] and Fermi surfaces in generic dimensions [10, 11, 12], subleading logarithmic terms for corner contributions to 2D critical systems [13, 14] and Goldstone modes in symmetry-breaking phases [15, 16, 17, 18], and the topological EE for non-local orders [19, 20, 21, 22, 23].
@@
However, EE is not a faithful mixed-state entanglement measurement due to its incompetence in distinguishing quantum entanglement from classical correlation. Thus, many entanglement measurements for mixed states have been proposed [24, 25], including the entanglement negativity [26, 27, 28, 29] (referred to as “negativity” henceforth for brevity), which was designed base on positive partial transpose criteria for the separability of density matrices [30, 31]. The evaluation of negativity hinges on the partial transpose of the given density matrix and can be carried out straightforwardly through basic matrix manipulations without invoking any optimization. Hence, negativity has been employed to examine entanglement in finite-temperature Gibbs states or tripartite ground states in various systems, spanning from one-dimensional conformal field theory [32, 33, 34, 35], bosonic systems [36, 37, 38, 39, 40], spin systems [41, 42, 43, 44, 45, 46, 47], to topologically ordered phases [48, 49, 50, 51, 52, 53].
@@
In the case of fermionic systems, the definition of partial transpose needs to be adjusted to accommodate the anticommuting statistical property. There exist two different proposals for fermionic partial transpose and corresponding fermionic negativity, as discussed in Refs. [54, 55] and Refs. [56, 57, 58] respectively. Despite being a “computable entanglement measurement”, fermionic negativity is only analytically tractable in free systems, especially at finite temperatures, and there have been studies based on both the former definition [59, 39, 60, 61] and the latter definition [56, 62, 63, 64, 65]. Therefore, it is desirable to design a quantum Monte Carlo (QMC) algorithm for large-scale simulation of interacting fermionic systems in an unbiased manner, which is the main goal of this letter. Throughout this paper, we adopt the definition in Refs. [56, 57] under which the partial transpose of a Gaussian state remains a Gaussian state. Additionally, instead of utilizing the originally proposed negativity which involves trace norm of partially transposed density matrices (PTDMs) [28], we consider Rényi negativity which involves moments of PTDMs, as done in several previous studies on other systems [32, 33, 45, 37, 40, 39, 47].
@@
In fact, our main result is more broadly applicable. We show that generic PTDMs can be written as a weighted sum of Gaussian states, representing free fermions coupled with auxiliary fields, similar to Grover’s pioneering work on reduced density matrices for EE [66]. Our finding facilitates the calculation of Rényi negativity in a tractable manner, thus establishing it as a powerful tool for characterizing entanglement in mixed states of interacting fermions. We demonstrate this relation using determinant quantum Monte Carlo (DQMC) simulations [67, 68, 69] on two paradigmatic models in the realm of strongly correlated electrons, namely, the Hubbard model and the spinless $t$-$V$ model. These two models on bipartite lattices at half-filling are sign-problem-free and both ground-state and finite-temperature properties can be feasibly simulated within the DQMC framework. The relation between negativity and finite temperature transition in fermionic systems is unveiled.
@@
Partially transposed density matrix in DQMC calculations.—Various definitions of negativity in the literature share a common and central dependency, namely, the partial transpose of the density matrix. In this work, we adopt the partial time-reversal transformation proposed by Shapourian et al. [56, 57] as the fermionic partial transpose.
@@
We begin with the general partitioning of a fermionic lattice model. It is defined using annihilation (creation) operators $c_{j\sigma}^{(\dagger)}$, which satisfy the anticommutation relations $\{c_{j\sigma},c^{\dagger}_{k\sigma^{\prime}}\}=\delta_{jk}\delta_{\sigma\sigma^{\prime}}$, where $j,k=1,\dots,N$ are the labels of the sites and $\sigma,\sigma^{\prime}$ are the indices for internal degrees of freedom such as spin. This lattice system, denoted as $A$, generally exists within a larger space. After tracing out the environment $\bar{A}$, system $A$ typically exists in a mixed state $\rho$. For example, if system $A$ is in contact with a much larger thermal bath at temperature $T$, then we obtain a finite-temperature Gibbs state $\rho=e^{-\beta H}/\mathrm{Tr}e^{-\beta H}$ with $\beta=1/T$ the inverse temperature and $H$ the Hamiltonian of the system $A$. Next, we further divide system $A$ into two parties belonging to two complementary spatial regions respectively, i.e., $A=A_{1}\cup A_{2}$. Then the density matrix acting on Hilbert space $\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ can be expanded as $\rho=\sum_{A_{1},A_{2},A_{1}^{\prime},A_{2}^{\prime}}\rho_{A_{1},A_{2};A_{1}^{\prime},A_{2}^{\prime}}|A_{1}\rangle|A_{2}\rangle\langle A_{1}^{\prime}|\langle A_{2}^{\prime}|$.
@@
The fermionic partial transpose of density matrix $\rho$ with respect to subsystem $A_{2}$, denoted as $\rho^{T_{2}^{f}}$, exhibits a highly succinct mathematical expression in the Majorana basis [56, 57, 70]. Under Majorana basis, an arbitrary density operator can be expressed as a constrained superposition of products of Majorana operators, which are defined as $\gamma_{2j-1,\sigma}=c_{j,\sigma}+c^{\dagger}_{j,\sigma}$ and $\gamma_{2j,\sigma}=-\mathrm{i}(c_{j,\sigma}-c_{j,\sigma}^{\dagger})$. It is found that $\rho^{T^{f}_{2}}$ can be obtained by applying the following transformation to the Majorana operators associated with subsystem $A_{2}$:
@@
Remarkably, under this definition, the fermionic partial transpose of a Gaussian state, denoted as $\rho_{0}\sim e^{\frac{1}{4}\bm{\gamma}^{T}W\bm{\gamma}}$, retains its Gaussian nature. The question then pertains to determining the explicit form of $\rho_{0}^{T_{2}^{f}}$ or $W^{T_{2}^{f}}$. To this end, it is important to emphasize that a Gaussian state $\rho_{0}$ can be alternatively characterized by the Green’s function $\Gamma_{kl}=\langle[\gamma_{k},\gamma_{l}]\rangle/2$, which is averaged with respect to $\rho_{0}$ itself and also called covariance matrix. This matrix is connected to the $W$ matrix through the relations $\tanh(-W/2)=\Gamma$ [54, 70]. By employing the definition of $\Gamma$ and the partial transpose in the Majorana basis (refer to Eq. (1)), the partial transpose of the covariance matrix can be formulated as
@@
where $\Gamma^{bb^{\prime}}$ ($b,b^{\prime}=1,2$) denotes the block comprising the matrix elements with rows pertaining to subsystem $A_{b}$ and columns pertaining to subsystem $A_{b^{\prime}}$. The Gaussian state described by $\Gamma^{T_{2}^{f}}$ precisely yields the $\rho_{0}^{T_{2}^{f}}$ through $\tanh(-{W^{T_{2}^{f}}}/{2})=\Gamma^{T_{2}^{f}}$  [^1] The proof uses the Wick theorem for Majorana monomial, see Ref. [54] and Supplementary Material (SM) [70] for details..
@@
It is now pertinent to redirect our attention towards the partial transpose for interacting fermionic systems, whose density matrices are not Gaussian states. Nonetheless, within the framework of DQMC, after Trotter decomposition and Hubbard-Stratonovich decoupling [70], the original two-particle interaction terms are replaced by fermion bilinears coupled with spacetime-dependent auxiliary fields $\mathbf{s}$. The partition function is given by $Z=\sum_{\mathbf{s}}\mathrm{Tr}[\prod_{l=1}^{L_{\tau}}e^{\mathbf{c}^{\dagger}K_{l}[\mathbf{s}]\mathbf{c}}]$, where $L_{\tau}$ is the number of time slices and $\mathbf{c}=(c_{1},\dots,c_{N})^{T}$ (for simplicity we abbreviate spin indices). The density matrix $\rho$ can also be expressed as a weighted sum of Gaussian operators, explicitly $\rho=\sum_{\mathbf{s}}P_{\mathbf{s}}\rho_{\mathbf{s}}$ with $P_{\mathbf{s}}$ representing the weight of configuration $\mathbf{s}$ [66, 70]. After partial transpose, it takes the form
@@
The aforementioned equations (4) and (5), along with Eq. (3), are the main result of this letter and can be employed to investigate negativity and negativity spectrum within the conventional DQMC framework, fully analogous to the analysis of EE and entanglement spectrum, respectively.
Rényi negativity for Hubbard chain.— We consider the half-filled Hubbard chain with periodic boundary conditions, described by the Hamiltonian
@@
which is a sign-problem-free model [69, 70]. We will benchmark DQMC results from two perspectives: (i) a numerical comparison with results obtained from exact diagonalization  [^2] We utilized the definition of fermionic partial transpose in Fock space for exact diagonalization, see the formula Eqs. (S2) and (S3) in SM [70], and (ii) providing a physical explanation for why the negativity is a more competent mixed-state entanglement measurement compared to EE in the context of a quantum-classical crossover [56, 62].
We define the rank-$n$ Rényi negativity as
@@
where the $n$-th moment of the PTDM, denoted as $\text{Tr}[(\rho^{T_{2}^{f}})^{n}]$  [^3] We note that in the context of studying fermionic systems, an alternative definition for the moments of the PTDM is given by $\text{Tr}(\rho^{T_{2}^{f}}\rho^{T_{2}^{f}\dagger}\cdots)$ [56, 62, 65], which aims to account for the non-Hermitian nature and potential complex eigenvalues associated with the partial time reversal of the density matrix. However, the models under consideration in this letter do not exhibit this issue due to specific symmetries they possess that circumvent the sign problem. , is also referred to as the replica approach of negativity in previous studies [32, 33]. The quantity $\mathcal{E}_{n}$ is formally a direct analog to rank-$n$ Rényi EE $S_{n}(A_{1})=-(\ln\mathrm{Tr}\rho^{n}_{A_{1}})/(n-1)$, where $\rho_{A_{1}}=\mathrm{Tr}_{A_{2}}\rho$ represents the reduced density operator obtained after tracing out subsystem $A_{2}$. Utilizing Eq. (4), we can derive the DQMC expression for measuring, for instance, the rank-two Rényi negativity,
$\mathcal{E}_{2}=$ $-\ln\Big{\{}\sum_{{\bf s}_{1}{\bf s}_{2}}P_{{\bf s}_{1}}P_{{\bf s}_{2}}$ (8)
×$\times\det\left[G_{\bf{s}_{1}}^{T_{2}^{f}}G_{\bf{s}_{2}}^{T_{2}^{f}}+\left(I-G_{\bf{s}_{1}}^{T_{2}^{f}}\right)\left(I-G_{\mathbf{s}_{2}}^{T_{2}^{f}}\right)\right]\Big{\}}$
@@
The distinction between the fermionic partial transpose and the conventional one is presented herein. For bosonic systems, considering that $\mathrm{Tr}[(\rho^{T_{2}})^{2}]=\mathrm{Tr}[\rho^{2}]$, the rank-two Rényi negativity becomes trivial, thereby rendering the minimal meaningful rank as three [32, 33, 47]. However, we show in the occupation number representation that all fermionic PTDMs satisfy ${\rm Tr}[(\rho^{T_{2}^{f}})^{2}]={\rm Tr}[(\rho\hat{X}_{2}(\pi))^{2}]$ with $\hat{X}_{2}\left(\theta\right)=e^{{\rm i}\theta\sum_{j\in A_{2}}{n}_{j}}$ being the disorder operator [70]. Consequently, $\mathcal{E}_{2}$ can reveal the entanglement information of the system. As shown in Fig. 1, the results calculated by DQMC and exact diagonalization show strong agreement in both the zero-temperature and finite-temperature regimes. In the zero-temperature regime, the pattern of the rank-two Rényi negativity exhibits analogous variations to those of the rank-two Rényi EE [66, 81], in response to alterations in the length of subsystem $A_{1}$, denoted as $L_{A_{1}}$. However, at finite temperatures, the negativity maintains a symmetric pattern, which is different from the behavior of EE [82, 81]. As the temperature rises, the magnitude of the negativity increases, resulting in an overall non-zero shift corresponding to a non-zero thermodynamic entropy of $-\ln(\mathrm{Tr}\rho^{2})$.
Figure 1: The variation of the rank-two Rényi negativity $\mathcal{E}_{2}$ for a six-site Hubbard chain with periodic boundary conditions is depicted as a function of the subsystem length $L_{A_{1}}$. The solid lines represent the exact diagonalization results, which agree with the DQMC results at both (a) zero temperature and (b) finite temperatures.
@@
Based on the above observation in finite-temperature regime, we also examine the ratio between $\text{Tr}[(\rho^{T_{2}^{f}})^{n}]$ and $\text{Tr}[\rho^{n}]$ as previous Monte Carlo investigations on bosonic systems [45, 37, 47]

### v4

@@ Abstract
Many-body entanglement unveils additional aspects of quantum matter and offers insights into strongly correlated physics. While ground-state entanglement has received much attention in the past decade, the study of mixed-state quantum entanglement using negativity in interacting fermionic systems remains largely unexplored. We demonstrate that the partially transposed density matrix of interacting fermions, similar to their reduced density matrix, can be expressed as a weighted sum of Gaussian states describing free fermions, enabling the calculation of rank-$n$ Rényi negativity within the determinant quantum Monte Carlo framework. We calculate the rank-two Rényi negativity for the half-filled Hubbard model and the spinless $t$-$V$ model. Our calculation reveals that the area law coefficient of the Rényi negativity for the spinless $t$-$V$ model has a logarithmic finite-size scaling at the finite-temperature transition point. Our work contributes to the calculation of entanglement and sets the stage for future studies on quantum entanglement in various fermionic many-body mixed states.
I Introduction
@@
The characterization of emerging quantum many-body phenomena is multifaceted. Traditionally, physicists have relied on local measurements based on linear response to investigate matter. In recent decades, the utilization of quantum entanglement, a fundamental concept in quantum physics and a powerful tool in quantum information, has become pivotal in unveiling the additional aspects of quantum matter, including the identification of exotic phases and quantum criticality [1, 2, 3]. A prominent example is the entanglement entropy (EE) used in bipartite ground-state entanglement studies [4], where various corrections to the area law [5] have been employed to classify quantum phases. These include logarithmic corrections in the leading area-law term for 1D critical systems [6] and Fermi surfaces in generic dimensions [7], subleading logarithmic terms for corner contributions to 2D critical systems [8] and Goldstone modes in symmetry-breaking phases [9], and the topological EE for non-local orders [10, 11].
@@
However, EE is not a faithful mixed-state entanglement measurement due to its incompetence in distinguishing quantum entanglement from classical correlation. Thus, many entanglement measurements for mixed states have been proposed [12], including the entanglement negativity [13, 14, 15, 16] (referred to as “negativity” henceforth for brevity), which was designed base on positive partial transpose criteria for the separability of density matrices [17, 18]. The evaluation of negativity hinges on the partial transpose of the given density matrix and can be carried out straightforwardly through basic matrix manipulations without invoking any optimization. Hence, negativity has been employed to examine entanglement in finite-temperature Gibbs states or tripartite ground states in various systems, spanning from one-dimensional conformal field theory [19, 20, 21], bosonic systems [22, 23, 24, 25], spin systems [26, 27, 28, 29, 30], to topologically ordered phases [31, 32, 33, 34, 35].
@@
In the case of fermionic systems, the definition of partial transpose needs to be adjusted to accommodate the anticommuting statistical property. There exist two different proposals for fermionic partial transpose (FPT) and corresponding fermionic negativity, as discussed in Refs. [36, 37] and Refs. [38, 39, 40] respectively. Despite being a computable entanglement measurement, fermionic negativity is only analytically tractable in free systems, especially at finite temperatures, and there have been studies based on both the former definition [41, 25, 42, 43] and the latter definition [38, 44, 45]. Therefore, it is desirable to design a quantum Monte Carlo (QMC) algorithm for large-scale simulation of interacting fermionic systems in an unbiased manner, which is the main goal of this work. Throughout this paper, we adopt the definition in Refs. [38, 39] under which the partial transpose of a Gaussian state remains a Gaussian state. Additionally, instead of utilizing the originally proposed negativity which involves trace norm of partially transposed density matrices (PTDMs) [15], we consider Rényi negativity (RN) which involves moments of PTDMs, as done in several previous studies on other systems [19, 20, 28, 23, 25, 30].
@@
In fact, our main result is more broadly applicable. We show that generic PTDMs can be written as a weighted sum of Gaussian states, representing free fermions coupled with auxiliary fields, similar to Grover’s pioneering work on reduced density matrices for EE [46]. Our finding facilitates the calculation of RN in a tractable manner, thus establishing it as a powerful tool for characterizing entanglement in mixed states of interacting fermions. We demonstrate this relation using determinant quantum Monte Carlo (DQMC) simulations [47, 48, 49] on two paradigmatic models in the realm of strongly correlated electrons, namely, the Hubbard model and the spinless $t$-$V$ model. These two models on bipartite lattices at half-filling are sign-problem-free and both ground-state and finite-temperature properties can be feasibly simulated within the DQMC framework. The relation between negativity and finite temperature transition in fermionic systems is unveiled.
II Results
II.1 Partially transposed density matrix integrated to DQMC framework
@@
Various definitions of negativity in the literature share a common and central dependency, namely, the partial transpose of the density matrix. In this work, we adopt the partial time-reversal transformation proposed by Shapourian et al. [38, 39] as the FPT.
@@
We begin with the general partitioning of a fermionic lattice model. It is defined using annihilation (creation) operators $c_{j\sigma}^{(\dagger)}$, which satisfy the anticommutation relations $\{c_{j\sigma},c^{\dagger}_{k\sigma^{\prime}}\}=\delta_{jk}\delta_{\sigma\sigma^{\prime}}$, where $j,k=1,\dots,N$ are the labels of the sites and $\sigma,\sigma^{\prime}$ are the indices for internal degrees of freedom such as spin. In the following discussion, we may use a column vector $\mathbf{c}=(c_{1,\uparrow},\dots,c_{N,\uparrow},c_{1,\downarrow},\dots,c_{N,\downarrow})^{T}$ to compactly encapsulate all the fermionic operators. This lattice system, denoted as $A$, generally exists within a larger space, as illustrated in Fig. 1A. After tracing out the environment $\bar{A}$, system $A$ typically exists in a mixed state $\rho$. For example, if system $A$ is in contact with a much larger thermal bath at temperature $T$, then we obtain a finite-temperature Gibbs state $\rho=e^{-\beta H}/\mathrm{Tr}e^{-\beta H}$ with $\beta=1/T$ the inverse temperature and $H$ the Hamiltonian of the system $A$. Next, we further divide system $A$ into two parties belonging to two complementary spatial regions respectively, i.e., $A=A_{1}\cup A_{2}$. Then the density matrix acting on Hilbert space $\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ can be expanded as $\rho=\sum_{A_{1},A_{2},A_{1}^{\prime},A_{2}^{\prime}}\rho_{A_{1},A_{2};A_{1}^{\prime},A_{2}^{\prime}}|A_{1}\rangle|A_{2}\rangle\langle A_{1}^{\prime}|\langle A_{2}^{\prime}|$.
@@
The FPT of density matrix $\rho$ with respect to subsystem $A_{2}$, denoted as $\rho^{T_{2}^{f}}$, exhibits a highly succinct mathematical expression in the Majorana basis [38, 39]. Under Majorana basis, an arbitrary density operator can be expressed as a constrained superposition of products of Majorana operators, which are defined as $\gamma_{2j-1,\sigma}=c_{j,\sigma}+c^{\dagger}_{j,\sigma}$ and $\gamma_{2j,\sigma}=-\mathrm{i}(c_{j,\sigma}-c_{j,\sigma}^{\dagger})$. It is found that $\rho^{T^{f}_{2}}$ can be obtained by applying the following transformation to the Majorana operators associated with subsystem $A_{2}$:
@@
Remarkably, under this definition, the partial transpose of a Gaussian state, denoted as $\rho_{0}\sim e^{\frac{1}{4}\bm{\gamma}^{T}W\bm{\gamma}}$ with $\bm{\gamma}=(\gamma_{1,\uparrow},\dots,\gamma_{2N,\uparrow},\gamma_{1,\downarrow},\dots,\gamma_{2N,\downarrow})^{T}$, retains its Gaussian nature. The question then pertains to determining the explicit form of $\rho_{0}^{T_{2}^{f}}$ or $W^{T_{2}^{f}}$. To this end, it is important to emphasize that a Gaussian state $\rho_{0}$ can be alternatively characterized by the Green’s function $\Gamma_{kl}=\langle[\gamma_{k},\gamma_{l}]\rangle/2$, which is averaged with respect to $\rho_{0}$ itself and also called covariance matrix. This matrix is connected to the $W$ matrix through the relation $\tanh(-W/2)=\Gamma$, or inversely, $W=\ln\left[(I+\Gamma)^{-1}(I-\Gamma)\right]$ [50] (see also the Supplementary Note 3 for a proof). By employing the definition of $\Gamma$ and the partial transpose in the Majorana basis (refer to Eq. (1)), the partial transpose of the covariance matrix can be formulated as
@@
where $\Gamma^{bb^{\prime}}$ ($b,b^{\prime}=1,2$) denotes the block comprising the matrix elements with rows pertaining to subsystem $A_{b}$ and columns pertaining to subsystem $A_{b^{\prime}}$. The Gaussian state described by $\Gamma^{T_{2}^{f}}$ precisely yields $\rho_{0}^{T_{2}^{f}}$ through the relation $\tanh(-{W^{T_{2}^{f}}}/{2})=\Gamma^{T_{2}^{f}}$, i.e., $(\rho_{0}[\Gamma])^{T_{2}^{f}}=\rho_{0}[\Gamma^{T^{f}_{2}}]$ with $\rho_{0}[\Gamma]\sim e^{\bm{\gamma}^{T}\ln\left[(I+\Gamma)^{-1}(I-\Gamma)\right]\bm{\gamma}}$. This elegant fact is proved using Wick’s theorem for Majorana monomials [36] (see the Supplementary Note 3 for details).
@@
Figure 1: A schematic illustration of the core concepts of this work. (a) Illustration of the general tripartite geometry. First, we trace out the environment $\bar{A}$ to obtain the reduced density matrix $\rho_{A}$, and then evaluate the entanglement between subsystems by either tracing out or partially transposing $A_{2}$. (b) The fermionic partial transpose of Gaussian states, $\rho_{0}[G]\sim e^{\mathbf{c}^{\dagger}\ln(G^{-1}-I)\mathbf{c}}$, remains Gaussian. These states represent free fermions, including the auxiliary-field-dependent density matrix $\rho_{\mathbf{s}}$ in the DQMC framework. The reduced and partially transposed density matrices thus share a unified expression in terms of the corresponding Green’s functions. Specifically, the reduced Green’s function is $G^{\rm red}_{\mathbf{s},A_{1}}=\langle c_{j}c_{k}^{\dagger}\rangle_{\mathbf{s}}$ for $j,k\in A_{1}$, while the partially transposed Green’s function $G_{\mathbf{s},A_{1}}^{\rm FPT}=G_{\mathbf{s}}^{T_{2}^{f}}$ is given by Eq. (3).
It is now pertinent to redirect our attention towards the partial transpose for interacting fermionic systems, whose density matrices are not Gaussian states. Nonetheless, within the framework of DQMC, the original Hamiltonian $H$ is transformed by replacing interaction terms with fermion bilinears coupled to spacetime-dependent auxiliary fields $\mathbf{s}$ [47, 48, 49]. Specifically, we consider the Gibbs state $\rho=e^{-\beta H}/Z$, where $H$ consists of a free-fermion term $H_{0}$ and a two-particle interaction term $H_{I}$. We employ Trotter decomposition to factorize the exponential operator $e^{-\beta H}$ as $(e^{-\Delta_{\tau}H})^{L_{\tau}}=\prod_{l=1}^{L_{\tau}}e^{-\Delta_{\tau}H_{I}}e^{-\Delta_{\tau}H_{0}}+O(\Delta_{\tau}^{2})$ with $L_{\tau}=\beta/\Delta_{\tau}$ being the number of imaginary-time slices. We then apply a Hubbard-Stratonovich (HS) transformation to decouple the interaction terms across different time slices. This procedure yields
$\rho=\frac{1}{Z}\sum_{\mathbf{s}}\eta\left[\mathbf{s}\right]\prod_{l=1}^{L_{\tau}}\left(e^{\mathbf{c}^{\dagger}V\left[\mathbf{s}(l)\right]\mathbf{c}}e^{-\Delta_{\tau}H_{0}}\right)\equiv\sum_{\mathbf{s}}P_{\mathbf{s}}\rho_{\mathbf{s}},$ (4)
where each $\mathbf{s}$-configuration is distributed over both the imaginary-time and spatial directions, contributing a scalar factor $\eta\left[\mathbf{s}\right]$ and a product of Gaussian operators. Here, both $V[\mathbf{s}]$ and $\eta[\mathbf{s}]$ are derived from $H_{I}$, and their forms depend on the specific interactions and HS decoupling channels. For detailed expressions related to the two models examined in this study, please refer to the Supplementary Note 2. Since the product of Gaussian states remains a Gaussian state up to a normalization factor [50], the interacting fermionic density matrix $\rho\sim e^{-\beta H}$ can ultimately be written as a weighted sum of Gaussian operators $\rho_{\mathbf{s}}$, with $P_{\mathbf{s}}$ denoting the configuration weight [46]. By leveraging the linearity of the partial transpose, we can first individually compute the FPT for each Gaussian state $\rho_{\mathbf{s}}$. We then sum these results, weighted by their respective probabilities $P_{\mathbf{s}}$, to obtain the FPT of the entire density matrix $\rho$:
@@
The aforementioned equations (5) and (6), along with Eq. (3), are the main results of this work and can be employed to investigate negativity and negativity spectrum within the conventional DQMC framework, fully analogous to the analysis of EE and entanglement spectrum (see Fig. 1B), respectively.
II.2 Quantum-classical crossover in Hubbard chain
We first consider the half-filled Hubbard chain with periodic boundary conditions, illustrated in Fig. 2A and described by the Hamiltonian
@@ II.2 Quantum-classical crossover in Hubbard chain
which is a sign-problem-free model [49]. We will benchmark DQMC results from two perspectives: (i) a numerical comparison with results obtained from exact diagonalization (ED) where we employ the definition of FPT in the Fock space (see Supplementary Eqs. (2) and (3)), and (ii) providing a physical explanation for why the negativity is a more competent mixed-state entanglement measurement compared to EE in the context of a quantum-classical crossover [38, 44].
We define the rank-$n$ RN as
@@
where the $n$-th moment of the PTDM, denoted as $\text{Tr}[(\rho^{T_{2}^{f}})^{n}]$ , is also referred to as the replica approach of negativity in previous studies [19, 20]. The quantity $\mathcal{E}_{n}$ is formally a direct analog to rank-$n$ Rényi EE $S_{n}(A_{1})=-(\ln\mathrm{Tr}\rho^{n}_{A_{1}})/(n-1)$, where $\rho_{A_{1}}=\mathrm{Tr}_{A_{2}}\rho$ represents the reduced density operator obtained after tracing out subsystem $A_{2}$. Utilizing Eq. (5), we can derive the DQMC expression for measuring, for instance, the rank-two RN
$\mathcal{E}_{2}=$ $-\ln\Big{\{}\sum_{{\bf s}_{1}{\bf s}_{2}}P_{{\bf s}_{1}}P_{{\bf s}_{2}}\det\left[G_{\bf{s}_{1}}^{T_{2}^{f}}G_{\bf{s}_{2}}^{T_{2}^{f}}+\left(I-G_{\bf{s}_{1}}^{T_{2}^{f}}\right)\left(I-G_{\mathbf{s}_{2}}^{T_{2}^{f}}\right)\right]\Big{\}}.$ (9)
@@
The distinction between the FPT and the conventional one is presented herein. For bosonic systems, considering that $\mathrm{Tr}[(\rho^{T_{2}})^{2}]=\mathrm{Tr}[\rho^{2}]$, the rank-two RN becomes trivial, thereby rendering the minimal meaningful rank as three [19, 20, 30]. However, we show in the occupation number representation that all fermionic PTDMs satisfy ${\rm Tr}[(\rho^{T_{2}^{f}})^{2}]={\rm Tr}[(\rho\hat{X}_{2}(\pi))^{2}]$ with $\hat{X}_{2}\left(\theta\right)=e^{{\rm i}\theta\sum_{j\in A_{2}}{n}_{j}}$ being the disorder operator (see the Supplementary Note 1 for details). Consequently, $\mathcal{E}_{2}$ can reveal the entanglement information of the system. As shown in Fig. 2B, the results calculated by DQMC and ED show strong agreement in both the zero-temperature and the finite-temperature regimes. In the former regime, the pattern of the rank-two RN exhibits analogous variations to those of the rank-two Rényi EE [46, 52], in response to alterations in the length of subsystem $A_{1}$, denoted as $L_{A_{1}}$. However, at finite temperatures, the negativity maintains a symmetric pattern, which is different from the behavior of EE [53, 52]. As the temperature rises, the magnitude of the negativity increases, resulting in an overall non-zero shift corresponding to a non-zero thermodynamic entropy of $-\ln(\mathrm{Tr}\rho^{2})$.
Figure 2: Rank-two RN(R) in the half-filled Hubbard chain. (a) A schematic illustration of the half-filled Hubbard chain with periodic boundary conditions and its bipartition geometry. (b) The variation of the rank-two RN $\mathcal{E}_{2}$ for a six-site Hubbard chain is depicted as a function of the subsystem length $L_{A_{1}}$. The solid lines represent the ED results, which agree with the DQMC results at both zero temperature (left panel) and finite temperatures (right panel). (c) Quantum-classical crossover. The scaled RNR $R_{2}/L$ and EE $S_{2}/L$ of the half-filled Hubbard chain under a half-chain bipartition (i.e., we take $L_{A_{1}}=L/2$) vary as functions of temperature. As the temperature rises, the scaled EE for different lengths increases and converges, indicating a dominance of volume law at high temperatures. Meanwhile, the RNR begins to vanish once the temperature reaches a critical value associated with the finite-size gap $1/L$ [44]. The error bars in (b) and (c) represent the standard errors from Monte Carlo sampling.
@@
Based on the above observation at finite temperatures, we also examine the ratio between $\text{Tr}[(\rho^{T_{2}^{f}})^{n}]$ and $\text{Tr}[\rho^{n}]$ dubbed the Rényi negativity ratio (RNR) [28, 23, 30]